# Randomized Black-Box PIT for Small Depth +-Regular Non-commutative Circuits

### v1

@@ 4.4 Coefficient Modification by Modulo Counting Automaton
Lemma 4.4.1.
Let $\tilde{f}=\hat{f}_{1}+F_{1}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ be the non-zero non-commutative polynomial obtained in Step 1. Let $\hat{f}^{(c)}_{3}$ be the commutative polynomial obtained from $\hat{f}_{1}$ after Steps (2) and (3) and let $F^{(c)}_{3}$ be the commutative polynomial obtained from $F_{1}$ after Steps (2) and (3). If $\hat{f}^{(c)}_{3}+F^{(c)}_{3}=0$, then it is possible to modify the coefficients of $\tilde{f}$ obtained in Step (1) before executing Steps (2) and (3) such that the resulting non-commutative polynomial $\tilde{f}^{{}^{\prime}}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ can be transformed into a commutative polynomial through Steps (2) and (3) while ensuring that the resulting commutative polynomial retains its non-zeroness.
@@ 4.5 Black-box Randomized PIT for $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ Circuits
It is important to note that, by the matrix composition lemma (see Lemma 3.0.1), the polynomial $f^{(c)}$ is equal to the polynomial obtained as the output of Step 3, which is $\mathbf{C}[1,s^{2}]$. This summation arises because the automaton for Step (2) has two accepting states, leading to two entries in the final composed matrix.
If $f^{(c)}\in\mathbb{F}[W]$ is identically zero, we can apply the coefficient modification step (See section 4.4) to obtain a non-zero commutative polynomial. The maximum dimension of the substitution matrix in this modification step is $O(s^{2})$ (see Figures 5 and 4), which will result in a final matrix of dimension $O(s^{6})$ due to the matrix composition lemma (Lemma 3.0.1).
@@
Let $f$ be a non-commutative polynomial of degree $D$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then, $f$ can be transformed into an $s^{2}$-ordered power-sum polynomial while preserving its non-zeroness.
@@ 5.1 Transforming $f$ into an ordered power-sum polynomial
To demonstrate Theorem 5, we first establish that the polynomial $f$ can be transformed into a more structured n.c. polynomial that facilitates the testing for non-zeroness. Specifically, we show that $f$ can be converted into a $c$-ordered power-sum polynomial while preserving its non-zeroness, where $c$ is a function of the depth $d$ and size $s$. We have the following theorem.
@@
Let $f$ be a non-commutative polynomial of degree $D$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $+$-regular circuit of size $s$ and depth $d$. We denote the number of addition (i.e., $\sum$) layers in the circuit by $d^{+}$.
@@
    Here, each $Q_{ij}$ (for $i\in[s],j\in[D_{2}]$) can be computed by a $+$-depth $(d^{+}-1)$ circuit of size at most $s$. We denote the degree of $Q_{ij}$ polynomials as $D_{1}$. Therefore, $f$ is a sum of products of polynomials computed by $+$-depth $(d^{+}-1)$ circuits of size at most $s$. By induction hypothesis, for all $i\in[s],j\in[D_{2}]$, each $Q_{ij}$ can be converted into an $s^{(d^{+}-2)}$-ordered power-sum polynomial $\hat{Q}_{ij}$ while preserving non-zeroness.
    Next, we need to establish that these individual conversions preserve the non-zeroness of the transformed polynomial. As noted in Remark 4.1.3, the proof of Lemma 4.1.1 can be generalized to affirm this case. Specifically, the n.c. polynomial $\hat{f}_{1}$ defined as
@@
    Following this, the polynomial $\hat{f}_{1}$ can be product sparsified in a manner analogous to the depth-5 case. As highlighted in Remark 4.2.1, the proof of Lemma 4.2.1 can be extended to product sparsify the polynomial $\hat{f}_{1}$. Due to the structured nature of $\hat{Q}_{ij}$ polynomials, boundaries can be identified by an automaton. Thus, we can construct an automaton $\mathcal{A}_{2}$ similar to the one in Figure B.1. We consider the output of $\mathcal{A}_{2}$ on $\hat{f}_{1}$ and we can replace the commutative variables from this resulting polynomial with scalars, yielding a new n.c. polynomial $\hat{f}_{2}$ while maintaining non-zeroness.
    Finally, the polynomial $\hat{f}_{2}$ can be converted into a commutative polynomial while maintaining non-zeroness, similar to the depth-5 case. As noted in Remark 4.3.1, the proof of Lemma 4.3.1 can be generalized to convert $\hat{f}_{2}$ into a commutative polynomial $\hat{f}_{3}$ that preserves non-zeroness. Similar to the depth-5 case, boundaries of $\hat{Q}_{ij}$ polynomials can be identified by an automaton due to their structured nature. Thus, we can construct an automaton $\mathcal{A}_{3}$ similar to one in Figure C.1. We consider the output of $\mathcal{A}_{3}$ on $\hat{f}_{2}$. The resulting commutative polynomial is denoted $\hat{f}_{3}$. The commutative polynomial $\hat{f}_{3}$ is an $s^{(d^{+}-1)}$-ordered power-sum polynomial.
@@
5.1.1 An Automaton for Theorem 6
Let $f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$ be a polynomial computed by a depth-$d$ circuit of size $s$. We denote the number of addition (i.e., $\sum$) layers in the circuit by $d^{+}$.
Suppose we have substitution matrices $\mathbf{M}=(\mathbf{M_{x_{1}},M_{x_{2}},\ldots,M_{x_{n}}})$ that transforms each $Q_{ij}$ polynomials into an $s^{(d^{+}-2)}$-ordered power-sum polynomial while preserving its non-zeroness.. Let $\mathcal{A}$ be the corresponding substitution automaton with these substitution matrices $\mathbf{M}$.
@@ 5.1.1 An Automaton for Theorem 6
as follows. In section 4.5.1, we described how to construct a substitution automaton that converts a polynomial $f$ computed by a depth-5 circuit into an $s^{2}$-ordered power-sum polynomial $f^{(c)}$ while preserving its non-zeroness. We can treat $f^{(c)}$ as a n.c. polynomial.
Given that $f$ is a sum of products of $Q_{ij}$ polynomials and that the boundaries of these $Q_{ij}$ polynomials are unknown, we can modify the substitution automaton $\mathcal{A}$ to guess the boundaries of each $Q_{ij}$. This involves adding transitions to the starting state of automaton $\mathcal{A}$ (similar to the automaton depicted in Figure 1). We denote the new substitution automaton by $\mathcal{A^{\prime}}$
@@
where $F_{1}$ represents the spurious part resulting from incorrect guesses made by the automaton $\mathcal{A}^{\prime}$. Here,
@@
  • Automaton size for coefficients modification step: The size of the substitution automaton that filters out one spurious monomial from the polynomial $F_{1}$ (obtained by Step 1) is at most $14s$. This is due to the prime number $p$ required for this step being bounded by $\leq 4.4s$ (see Lemma D.0.1). Additionally, we need to 3 states to process the $\xi$-patterns (see Definition 4.0.2). This automaton construction is similar to the automaton illustrated in Figure 4 in the depth-5 case.
    The size of the substitution automaton that modifies the coefficients of the polynomial $\hat{f}_{1}+F_{1}$ (also obtained by Step 1) is bounded by $34.4s^{(d^{+}-1)}+$ 3. This is analogous to Figure 5 in the depth-5 case. Each $\xi$-pattern (see Definition 4.0.2) of $\hat{f}_{1}$ contains $s^{(d^{+}-2)}$ variables, and the prime number $p$ required for this step is bounded by $4.4\log D$, where $D$ is the degree of the $f$ and bounded by $2^{s}$. Therefore, with one starting state and two additional accepting states, the automaton size becomes $34.4s^{(d^{+}-1)}+$ 3.
    Consequently, the maximum size of the automaton for this step is $s^{(d^{+}-1)}+3$.

### v2

@@ 4.4 Coefficient Modification by Modulo Counting Automaton
Lemma 4.4.1 (Coefficient Modification Lemma).
Let $\tilde{f}=\hat{f}_{1}+F_{1}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ be the non-zero non-commutative polynomial obtained in Step 1. Let $\hat{f}^{(c)}_{3}$ and $F^{(c)}_{3}$ be the commutative polynomials obtained from $\hat{f}_{1}$ and $F_{1}$ after Steps (2) and (3), respectively. If $\hat{f}^{(c)}_{3}+F^{(c)}_{3}=0$, then the coefficients of $\tilde{f}$ can be modified before Steps (2) and (3) in such a way that the resulting non-commutative polynomial $\tilde{f}^{{}^{\prime}}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ (after coefficients modification step) can be transformed into a commutative polynomial through Steps (2) and (3) while ensuring that the resulting commutative polynomial retains its non-zeroness.
In particular, there exists a small substitution automaton of size $O(s^{2})$ that performs this commutative transformation.
In other words, if $\hat{f}^{(c)}_{3}+F^{(c)}_{3}=0$, then for each non-commutative variable $\xi_{i}\in[s]$, there exists an $O(s^{2})$-dimensional matrix substitution—acting as a transition matrix of a substitution automaton— such that when the n.c. polynomial $\tilde{f}$ (obtained after Step 1) is evaluated on these matrices, the coefficients of $\tilde{f}$ are modified, resulting in a non-commutative polynomial $\tilde{f}^{{}^{\prime}}\in\mathbb{F}\mathopen{\langle}\xi\mathclose{\rangle}$ that maintains its non-zero property. Furthermore, this newly obtained non-commutative polynomial $\tilde{f}^{{}^{\prime}}$ can be transformed into a commutative polynomial through Steps (2) and (3), while ensuring that the resulting commutative polynomial retains its non-zero property.
@@ 4.5 Black-box Randomized PIT for $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ Circuits
This differs from previous works ([AJMR19, BW05]) where only a single matrix entry was used as the output. It is important to note that, by the matrix composition lemma (see Lemma 3.0.1), the polynomial $f^{(c)}$ is equal to the polynomial obtained as the output of Step 3, which is $\mathbf{C}[1,s^{2}]$. This summation arises because the automaton for Step (2) has two accepting states, leading to two entries in the final composed matrix.
If $f^{(c)}\in\mathbb{F}[W]$ is identically zero, we can apply the coefficient modification step (See section 4.4) to obtain a non-zero commutative polynomial. The maximum dimension of the substitution matrix in this coefficient modification step is $O(s^{2})$ (see Figures 5 and 4), which will result in a final matrix of dimension $O(s^{6})$ due to the matrix composition lemma (Lemma 3.0.1).
@@
Let $f$ be a non-zero non-commutative polynomial of degree $D$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then, $f$ can be transformed into an $s^{2}$-ordered power-sum polynomial while preserving its non-zeroness. In particular, there exists a small substitution automaton of size $O(s^{6})$ that performs this transformation.
In other words, for each n.c. input variable $x_{i}$ (where $i\in[n]$) in $f$, there exists an $O(s^{6})$-dimensional matrix—acting as a transition matrix of a substitution automaton of size $O(s^{6})$. By evaluating $f$ on these matrices, the polynomial $f$ is transformed into an $s^{2}$-ordered power-sum polynomial while preserving its non-zeroness (in one particular entry of the resulting matrix).
@@ 5 Black-Box Randomized PIT for Small Depth +-Regular Circuits
Let $f$ be a non-commutative polynomial of degree $D$ over $X=\{x_{1},\ldots,x_{n}\}$, computed by a $+$-regular circuit of size $s$ and depth $d$. Then, $f\not\equiv 0$ if and only if $f$ is not identically zero on $\mathbb{M}_{N}(\mathbb{F})$, where $N=s^{O(d^{2})}$ and $|\mathbb{F}|$ is sufficiently large.
@@ 5.1 Transforming $f$ into an ordered power-sum polynomial
To demonstrate Theorem 5, we first establish that the polynomial $f$ can be transformed into a more structured n.c. polynomial that facilitates the testing for non-zeroness. Specifically, we show that $f$ can be converted into a $c$-ordered power-sum polynomial while preserving its non-zeroness, where $c$ is a function of the depth $d$ and size $s$. To prove this conversion, we need a generalization of Lemma 4.1.1; For completeness, we state this generalized lemma (see Lemma 5.1.1). Its proof follows a similar structure to that of Lemma 4.1.1. We note this as a remark here.
Remark 5.1.1.
The proof of Lemma 4.1.1 is based on the observation that if each polynomial $Q_{ij}$ (where $i\in[s]$ and $j\in[D_{2}]$, computable by a depth-3 +-regular circuit) can be converted into a n.c. polynomial $\hat{Q}_{ij}$ while preserving non-zeroness, then this lemma asserts that this transformation can be extended to the entire polynomial computed by the depth-5 circuit. Specifically, each $Q_{ij}$ can be converted into the n.c. polynomial $\hat{Q}_{ij}$, leading to the creation of the n.c. polynomial $\hat{f}$, while maintaining the non-zeroness of the entire polynomial. This can be extended to higher depths as follows:
Suppose
  1. 1.
    $f\in\mathbb{F}\langle x_{1},\ldots,x_{n}\rangle$ is computed by the sum of the products of $Q_{ij}$ polynomials, that is
    $f=\sum\limits_{i\in[s]}\prod\limits_{j\in[D_{2}]}Q_{ij}$
    where the degrees of $Q_{ij}$ are the same for all $i\in[s]$ and $j\in[D_{2}]$ and are computed by depth ($d-2$) $+$-regular circuits of size at most $s$.
  2. 2.
    Each $Q_{ij}$ individually can be converted into a n.c. polynomial $\hat{Q}_{ij}$ (which is an ordered power-sum polynomial) while preserving its non-zeroness. That is, $Q_{ij}=0$ if and only if $\hat{Q}_{ij}=0$.
Then this transformation can be extended to the entire polynomial $f$ computed by the depth-$d$ circuit as follows: each $Q_{ij}$ in $f$ can be individually converted to $\hat{Q}_{ij}$ to create the n.c. polynomial $\hat{f}$ from $f$, maintaining the non-zeroness of the entire polynomial.
The proof of this Lemma 5.1.1 can be found in Appendix A.1
Lemma 5.1.1.
Let
$f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$
be a non-commutative polynomial over $X=\{x_{1},\cdots,x_{n}\}$, computed by a depth $d$ $+$-regular circuit of size $s$, where each $Q_{ij}$ is computed by a depth $(d-2)$ $+$-regular circuit of size at most $s$. Let $D_{1}\geq s-1$ denote the degree of the polynomial $Q_{ij}$, for $i\in[s],j\in[D_{2}]$.
Suppose that any non-commutative polynomial $P$, computed by a depth $(d-2)$ $+$-regular circuit, can be converted into a non-commutative polynomial $\hat{P}$ (which is an ordered power-sum polynomial) while preserving non-zeroness. Then, each $Q_{ij}$ polynomial in $f$ can be converted to $\hat{Q}_{ij}$ to create the n.c. polynomial $\hat{f}=\sum_{i\in[s]}\prod_{j\in[D_{2}]}\hat{Q}_{ij}$. Then, if $f\neq 0$ then $\hat{f}\neq 0$ maintaining the non-zeroness of the entire polynomial $\hat{f}$. That is, $f=0$ if and only if $\hat{f}=0$.
Remark 5.1.2.
Unlike in Lemma 4.1.1, the $\hat{Q}_{ij}$ polynomials in Lemma 5.1.1 are not necessarily homogeneous. One of the hypotheses of the lemma states that any polynomial $P$, computed by a depth-$(d-2)$ $+$-regular circuit, can be transformed into an ordered power-sum polynomial $\hat{P}$ while preserving non-zeroness. However, $\hat{P}$ itself may be non-homogeneous. Importantly, the proof of Lemma 5.1.1 does not depend on whether $\hat{P}$ is homogeneous.
We will use this generalization to transform polynomials computed by larger depth $+$-regular circuits. It is important to note that, like Lemma 4.1.1, Lemma 5.1.1 is primarily aimed at establishing the non-zeroness of $\hat{f}$. However, in our approach, we cannot use a small automaton to directly transform $f$ into $\hat{f}$ without introducing a spurious part, $F$, as we did in the depth-5 case. In other words, when we apply a substitution automaton to transform $f$ the result is $\hat{f}+F$, where $F$ represents the spurious part.
We have the following theorem, which asserts that any non-zero polynomial $f$, computed by a depth-$d$ circuit, can be transformed into an ordered power-sum polynomial. The size of the automaton that performs this transformation is analyzed in section 5.1.1.
@@
Let $f$ be a non-zero non-commutative polynomial of degree $D$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $+$-regular circuit of size $s$ and depth $d$. Let $d^{+}$ be the number of addition (i.e., $\sum$) layers in the circuit.
@@
    Here, each $Q_{ij}$ (for $i\in[s],j\in[D_{2}]$) can be computed by a $+$-depth $(d^{+}-1)$ circuit of size at most $s$. That is, each $Q_{ij}$ can be computed by a depth-($d-2$) +-regular circuit.
    We denote the degree of $Q_{ij}$ polynomials as $D_{1}$. Therefore, $f$ is a sum of products of polynomials computed by $+$-depth $(d^{+}-1)$ circuits of size at most $s$. By induction hypothesis, for all $i\in[s],j\in[D_{2}]$, each $Q_{ij}$ can be converted into an $s^{(d^{+}-2)}$-ordered power-sum polynomial $\hat{Q}_{ij}$ while preserving non-zeroness. As noted in Remark 5.1.2, $\hat{Q}_{ij}$ polynomials can be non-homogenous in general.
    Next, we need to establish that these individual conversions preserve the non-zeroness of the transformed polynomial. By Lemma 5.1.1, this can be established. Specifically, the n.c. polynomial $\hat{f}_{1}$ defined as
@@
    Following this, the polynomial $\hat{f}_{1}$ can be product sparsified in a manner analogous to the depth-5 case. As $\hat{f}_{1}$ polynomial is represented as a sum of $s$ number of products of ordered power-sum polynomials, thus Lemma 4.2.1 can be used to product sparsify the polynomial $\hat{f}_{1}$. Due to the structured nature of $\hat{Q}_{ij}$ polynomials (ordered power-sum polynomials), boundaries can be identified by a small automaton. Thus, we can construct an automaton $\mathcal{A}_{2}$ similar to the one in Figure B.1. We consider the output of $\mathcal{A}_{2}$ on $\hat{f}_{1}$ and we can replace the commutative variables from this resulting polynomial with scalars, yielding a new n.c. polynomial $\hat{f}_{2}$ while maintaining non-zeroness.
    Finally, the polynomial $\hat{f}_{2}$ can be converted into a commutative polynomial while maintaining non-zeroness, similar to the depth-5 case. As $\hat{f}_{2}$ is represented as a sum of products of a small number of ordered power-sum polynomials, thus Lemma 4.3.1 can be used to convert $\hat{f}_{2}$ into a commutative polynomial $\hat{f}_{3}$ that preserves non-zeroness.
    Boundaries of $\hat{Q}_{ij}$ polynomials can be identified by a small automaton due to their structured nature. Thus, we can construct an automaton $\mathcal{A}_{3}$ similar to one in Figure C.1. We consider the output of $\mathcal{A}_{3}$ on $\hat{f}_{2}$. The resulting commutative polynomial is denoted $\hat{f}_{3}$. The commutative polynomial $\hat{f}_{3}$ is an $s^{(d^{+}-1)}$-ordered power-sum polynomial.
@@
5.1.1 Analyzing the Automaton Size for Theorem 6
Now we analyze the size of the automaton that transforms the given polynomial $f$ into an ordered power-sum polynomial in Theorem 6.
Before we proceed further, we recall that the result of Step 1 transformation in depth-5 case is a homogeneous n.c. polynomial $\tilde{f}=\hat{f}_{1}+F_{1}$. As noted the main advantages of Step 1 are
  1. 1.
    $\hat{f}_{1}$ is an ordered power-sum polynomial
  2. 2.
    boundaries of $\hat{Q}_{ij}$ polynomials (that is boundaries of monomials belonging to $\hat{Q}_{ij}$) can be efficiently identified using an automaton
However, after the product sparsification step (Step 2), the resulting polynomial $\hat{f}_{2}+F_{2}$ could become non-homogeneous. This is because the spurious part $F_{1}$ from Step 1 is a sum of the product of $\xi$-patterns (see Definition 4.0.2), and we cannot guarantee anything about the degree of each $\xi$-pattern in the monomials of $F_{1}$. But in each monomial in $\hat{f}_{1}$, each $\xi$-pattern has the same degree. In each monomial of the polynomial $\tilde{f}=\hat{f}_{1}+F_{1}$, the product sparsification step (Step 2) just selects some number of $\xi$-patterns as non-commutative, and the rest of the $\xi$-patterns as commutative. Therefore, the resulting polynomial $\hat{f}_{2}+F_{2}$ can be non-homogeneous in general (after removing commutative variables by scalar substitutions).
In Step 3, we simply convert the polynomial $\hat{f}_{2}+F_{2}$ into an ordered power-sum polynomial using fresh variables. These variables can be considered as non-commutative for subsequent applications using induction. Thus, the resulting ordered power-sum polynomial $\hat{f}_{3}+F_{3}$ can be non-homogeneous in general. Moreover, the coefficient modification step only updates the coefficient of the monomial (without changing any variables in it), thus it will not affect the degree.
The n.c. polynomial $\hat{f}_{2}+F_{2}$ obtained after the Step 2 transformation in the depth-5 case can be expressed as follows. Let $K$ represent the degree of $\hat{f}_{2}$. After the product sparsification step, the n.c. degree becomes $K=s\times D_{1}$, where $s$ is the number of terms in the sum and $D_{1}$ is the degree of the individual polynomials $Q_{ij}$.
We can write $\hat{f}_{2}+F_{2}$ as:
$\hat{f}_{2}+F_{2}=\left(\hat{f}_{2}+F^{K}_{2}\right)+F^{\neq K}_{2}$ (10)
where:
  • $F^{K}_{2}$ represents the sum of all monomials in $F_{2}$ that have degree exactly $K$.
  • $F^{\neq K}_{2}$ includes all the monomials in $F_{2}$ that have a degree different from $K$.
Now, observe that $\hat{f}_{2}$ and $F^{\neq K}_{2}$ are distinct, meaning their monomials do not interfere with each other’s coefficients. This is because the degree of the monomial does not match. This distinction remains true even after Step 3, as the transformation from Step 2 is bijective.
On the other hand, the monomials in $F^{K}_{2}$ could potentially affect the coefficients of monomials in $\hat{f}_{2}$. Specifically, if $\hat{f}_{2}+F^{K}_{2}=0$, this will hold true even after Step 3. To preserve non-zeroness, we can use the coefficient modification step (after Step 1), which ensures that $\hat{f}_{2}+F^{K}_{2}$ maintains non-zeroness throughout the process.
For higher depths, we need to redefine the notion of ”good” and ”spurious” parts. This is because we use only the degree to distinguish between the good and spurious parts. For example, consider the polynomial computed by a depth-7 circuit, which can be expressed as:
$g=\sum_{i\in[s]}\prod_{j\in[D_{3}]}P_{ij}$
where $P_{ij}$ is a polynomial computed by a depth-5 circuit. Let $f$ be $P_{ij}$ for some $i\in[s]$ and $j\in[D_{3}]$.
To apply the results from the depth-5 case to depth-7 circuits, we define the ”good” part of the polynomial $P_{ij}$ as $\hat{P}_{ij}=\hat{f}_{2}+F^{K}_{2}$. This is because we use only the degree to distinguish between the good and spurious parts, the monomials in $F^{K}_{2}$ are included in the good part for applying these results to higher depths. Specifically, the ”spurious” part of the Step 1 transformation of the polynomial computed by the depth-7 circuit includes all monomials whose degrees are not equal to $K$. Similar to the depth-5 case, in both the good and spurious parts each monomial after Step 1 is a product of ordered power-sum monomials.
Let
$f=\sum_{i\in[s]}\prod_{j\in[D_{2}]}Q_{ij}$
be a polynomial computed by a depth-$d$ circuit of size $s$. We denote the number of addition (i.e., $\sum$) layers in the circuit by $d^{+}$. In particular, $Q_{ij}$ can be computed by a depth ($d-2$) circuit.
Suppose we have substitution matrices $\mathbf{M}=(\mathbf{M_{x_{1}},M_{x_{2}},\ldots,M_{x_{n}}})$ that transforms each $Q_{ij}$ polynomials into an $s^{(d^{+}-2)}$-ordered power-sum polynomial while preserving its non-zeroness. We define that this $s^{(d^{+}-2)}$-ordered power-sum polynomial includes only the monomials where the degrees match. Any remaining monomials, which were obtained using a depth $(d-2)$ circuit, will be included in the spurious part $F_{1}$ of the depth-$d$ polynomial $f$.
Let $\mathcal{A}$ be the corresponding substitution automaton with these substitution matrices $\mathbf{M}$.
@@ 5.1.1 Analyzing the Automaton Size for Theorem 6
as follows. In section 4.5.1, we described how to construct a substitution automaton that converts a polynomial $f$ computed by a depth-5 circuit into an $s^{2}$-ordered power-sum polynomial $f^{(c)}$ while preserving its non-zeroness. We can treat $f^{(c)}$ as a n.c. polynomial by treating variables introduced in Step 3 as non-commutative.
Given that $f$ is a sum of products of $Q_{ij}$ polynomials and that the boundaries of these $Q_{ij}$ polynomials are unknown, we can modify the substitution automaton $\mathcal{A}$ to guess the boundaries of each $Q_{ij}$. This involves adding transitions to the starting state of automaton $\mathcal{A}$ (similar to the automaton depicted in Figure 1). For example, we need to add transitions from each accepting state of the automaton $\mathcal{A}$ to its starting state. We denote the new substitution automaton by $\mathcal{A^{\prime}}$
@@
where $F_{1}$ includes both the spurious sub-monomials resulting from incorrect boundary guesses made by the automaton $\mathcal{A}^{\prime}$ for $Q_{ij}$, as well as the ordered power-sum monomials (see Definition 4.0.2) produced by $\mathcal{A}^{\prime}$ where the degrees do not match (see, for example, Equation 10). Since the degrees of these sub-monomials do not match those of $\hat{Q}_{ij}$, any monomials containing these sub-monomials will be part of $F_{1}$. Thus, we can use claims similar to Claims 4.1.2, 4.1.1 to distinguish between the good and spurious parts.
Here,
@@
  • Automaton size for coefficients modification step: The size of the substitution automaton that filters out one spurious monomial from the polynomial $F_{1}$ (obtained by Step 1) is at most $14s$. This is due to the prime number $p$ required for this step being bounded by $\leq 4.4s$ (see Lemma D.0.1). Additionally, we need 3 states to process the $\xi$-patterns (ordered power-sum monomial, see Definition 4.0.2). This automaton construction is similar to the automaton illustrated in Figure 4 in the depth-5 case. We refer to each ordered power-sum monomial as a $\xi$-pattern, where $\xi$ denotes the set of variables used in the definition of an ordered power-sum polynomial (see Definition 4.0.1).
    The size of the substitution automaton that modifies the coefficients of the polynomial $\hat{f}_{1}+F_{1}$ (also obtained by Step 1) is bounded by $34.4s^{(d^{+}-1)}+$ 3. This is analogous to Figure 5 in the depth-5 case. Each $\xi$-pattern (see Definition 4.0.2) of $\hat{f}_{1}$ contains $s^{(d^{+}-2)}$ variables, and the prime number $p$ required for this step is bounded by $4.4\log D$, where $D$ is the degree of $f$ and bounded by $2^{s}$. Therefore, with one starting state and two additional accepting states, the automaton size becomes $34.4s^{(d^{+}-1)}+$ 3.
    Consequently, the maximum size of the automaton for this step is $34.4s^{(d^{+}-1)}+$ 3.